# Uncertainty-Guided Optimization on Large Language Model Search Trees

### v1

@@ Abstract
Beam search is a standard tree search algorithm when it comes to finding sequences of maximum likelihood, for example, in the decoding processes of large language models. However, it is myopic since it does not take the whole path from the root to a leaf into account. Moreover, it is agnostic to prior knowledge available about the process: For example, it does not consider that the objective being maximized is a likelihood and thereby has specific properties, like being bound in the unit interval. Taking a probabilistic approach, we define a prior belief over the LLMs’ transition probabilities and obtain a posterior belief over the most promising paths in each iteration. These beliefs are helpful to define a non-myopic Bayesian-optimization-like acquisition function that allows for a more data-efficient exploration scheme than standard beam search. We discuss how to select the prior and demonstrate in on- and off-model experiments with recent large language models, including Llama-2-7b, that our method achieves higher efficiency than beam search: Our method achieves the same or a higher likelihood while expanding fewer nodes than beam search.
@@ 1 Introduction
Beam search [18] is an optimization algorithm commonly applied to graph-, and in particular tree-structured problems. The number of possible paths that need to be considered in such settings typically grows exponentially, often exceeding the computational budget required to examine them all. This inevitably leads to computational uncertainty [12]: an uncertainty that could be fully resolved if enough compute was available to examine all paths, but in practice is present due to the limited resources. The standard beam search algorithm, while ubiquitous—e.g., in natural language processing for generating sentences under a large language model (LLM) and especially common in summarization and translation tasks [34, 46, 44, 37]—completely ignores this uncertainty.
@@
We posit that quantifying this uncertainty can be beneficial to ensure better explorations on LLMs’ search trees. To this end, we incorporate computational uncertainty into the search process to guide it in a non-myopic fashion (i.e., accounting for the belief about future nodes) and importantly, in a more data-efficient manner, akin to Bayesian optimization methods [BO, 21, 28, 6]. BO methods are recognized for their data efficiency, not merely because they quantify uncertainty, but because they exploit the structural characteristics within that uncertainty. E.g., in continuous optimization problems, prior knowledge, such as the smoothness of a function, is often available through Gaussian processes [30] or (Bayesian) neural networks [14, 19, 20].
@@
Figure 1: The intuition of our method—faded color represents unexplored subtrees. We sequentially expand a node of the LLM search tree based on the posterior belief. The latter is obtained by conditioning the unknown (the optimal total likelihood $\max c_{x_{0}\to x_{i}\to x_{n}}$ from the root $\boldsymbol{x}_{0}$ to a leaf $x_{n}$ when continuing down a particular node $x_{i}$) on the current observations (the total likelihood so far $c_{\boldsymbol{x}_{0}\to x_{i}}$ up until that node). This belief is induced by a simple prior belief on the LLM softmax probabilities $\boldsymbol{c}$. The samples of the posterior belief are used to score the unexplored subtrees.
@@
Unlike standard BO, however, beam search performs discrete optimization—its search space is a tree. Moreover, commonly, the values or rewards associated with each node are probabilities, bounded between 00 and $1$. Hence, uncertainty quantification here is not as well-studied as in standard BO. We will therefore assume that rewards at a node of the search tree are components of a discrete distribution. The characteristic property we aim to exploit is the concentration strength: whether the Categorical distributions are all highly concentrated at a few realistic options, or if some of them are nearly uniform, making any individual option drastically less likely to be the optimal one. Intuitively, one would expect this to have a strong influence on the number of paths that need to be considered. For instance, when the distribution is concentrated, it is less likely that other paths will overtake later on and one can be more greedy. Meanwhile, when the distribution has high entropy, the uncertainty about which token is best is higher and thereby requires more exploration and computational budget.
@@
In this work we are therefore interested in whether a probabilistic model that captures this aspect of the search space can help to decide which paths should be pursued and which can be ignored—see Fig. 1. Experiments on real-world text-generation benchmarks with GPT-2 [29] and Llama-2-7b [33] suggest that this is the case: Our method, called Uncertainty-guided Likelihood-Tree Search (ULTS), finds sentences with higher rewards than beam search with significantly fewer node expansions. ULTS only adds a small runtime overhead relative to the forward passes of the LLM.
In summary, our contribution is four-fold:
@@
    We propose ULTS: a probabilistic model on LLM search trees, where a prior belief is put on the LLM softmax outputs.
@@
    Based on that prior, we show how to easily sample from the implied posterior over optimal values, and how to use these samples to make decisions when searching the tree.
@@
    We demonstrate the efficiency of ULTS in decoding recent LLMs such as Llama-2-7b.
  4. (iv)
    We provide an open-source implementation at https://github.com/JuliaGrosse/ults, compatible with the transformers library [42].
@@ 2 Background
Let ${\mathcal{A}}=\{a_{i},\dots,a_{b}\}$ be a vocabulary—a set of natural-language tokens. A large language model (LLM) is a neural network that utilizes the attention mechanism [34], taking a context, an ordered sequence of tokens $\boldsymbol{x}_{0}=(x_{01},\dots,x_{0k})$—each component takes values in the set ${\mathcal{A}}$—and outputs a discrete probability distribution over ${\mathcal{A}}$, i.e. $\boldsymbol{c}=(c_{1},\dots,c_{b})$ where each $c_{i}\geq 0$ and $\smash{\sum_{i=1}^{b}}c_{i}=1$.
@@ 2.1 Decoding large language models
LLM decoding can be seen as a tree-search problem (Fig. 1). The context $\boldsymbol{x}_{0}$ is the root of the tree and at each step $i=1,\dots,d$, we are given a choice of $b$ many tokens to pick. This process is done recursively until termination; either when a specified depth $d$ has been reached or when a special token like “<EOS>” is selected. This means, the LLM search tree is an exponentially large tree (w.r.t. the number of tokens $b$ in the vocabulary) with depth $d$—the number of paths $\boldsymbol{x}_{0}\to x_{d}$ from the root to leaves is $b^{d}$. Considering the fact that $b$ ranges from around $32$k to $256$k [29, 3], even generating a short sequence of $d=50$ tokens requires search over an intractably large space.
@@
To address this problem, one can use a cheaper but heuristic way to explore this tree. The simplest way to do so is by sampling a token from the distribution $\boldsymbol{c}$ at each level of the tree [15]. In some natural language processing domains like machine translation and summarization, however, one often wants to find the sequence that maximizes the total/joint likelihood [^1] Implementation wise, this is usually done in the logarithmic space. $c_{\boldsymbol{x}_{0}\to x_{d}}:=\smash{\prod_{i=1}^{d}}c_{(i-1)i}$ associated with each possible sequence $(\boldsymbol{x}_{0},x_{1},\dots,x_{d})$ [40]. This is an optimization-on-tree problem—heuristic methods like greedy and beam search along with their variants [35, 25, 5, etc.] are the standard.
2.2 Sequential decision-making under uncertainty
The LLM decoding process described above is sequential—each $x_{i}$ is picked based on the previously seen sub-sequence $\boldsymbol{x}_{\text{prev}}=(\boldsymbol{x}_{0},x_{1},\dots,x_{i-1})$, without knowing the rest of the sequence $\boldsymbol{x}_{\text{next}}=(x_{i+1},\dots,x_{d})$. We must make a decision—selecting $x_{i}\in{\mathcal{A}}$—based on the observation $\boldsymbol{x}_{\text{prev}}$ so far, under the uncertainty surrounding the rest of the sequence $\boldsymbol{x}_{\text{next}}$. Finding a good path through the LLM search tree is thus an instance of sequential decision-making under uncertainty [17].
As in Bayesian optimization [21, 28], this can be approached by modeling a (probabilistic) belief about the unknown (in a tree-search problem, $\boldsymbol{x}_{\text{next}}$ or quantities that depend on it) given the current observation ($\boldsymbol{x}_{\text{prev}}$, etc.). This belief can then be used to make decision, by computing an acquisition function which scores possible realizations of $x_{i}$ given the belief above. The action that has the highest acquisition value is then selected and the process is repeated until termination.
2.3 Sequential decision-making on trees
@@
One can also perform a sequential decision-making under uncertainty on trees. First, notice that on each of the tree’s node $x_{i}$, we can define an optimal value $v_{x_{i}}$ associated with it: If $x_{d}$ is a leaf node, then $v_{x_{d}}$ corresponds to the total likelihood $c_{\boldsymbol{x}_{0}\rightarrow x_{d}}$ from the root until that leaf node. Meanwhile, for inner nodes $x_{i}$ it is defined recursively in a bottom-up fashion as the maximum of the children’s optimal values $\max_{x_{c}\in\mathrm{children}{(}x_{i})}{v_{x_{c}}}$. The optimization-on-tree problem can then be recast as finding a path $\boldsymbol{x}_{0}\to x_{d}$ that corresponds to the optimal value $v^{*}$ of the root.
@@
In this setting, the approach to perform sequential decision-making under uncertainty is then to perform probabilistic modeling over these values and use the posterior beliefs to make a decision about which path should be followed next. This kind of optimization has been done for game (e.g., Go) trees and directed-acyclic graphs [11, 8]. They assume that the optimal value $v_{x_{i}}$ of a node $x_{i}$ is decomposed into a latent score $g_{i}$ for the utility of the node if all remaining steps are taken randomly and additional increment $\Delta_{i}$ to quantify how much better this score can get when all remaining steps are taken optimally instead of randomly, i.e. $v_{x_{i}}=g_{i}+\Delta_{i}$. They assume a Gaussian process with the Brownian motion kernel as their prior belief.
@@ 3 ULTS: Uncertainty-guided Likelihood-Tree Search
In this section, we discuss our method: Uncertainty-guided Likelihood-Tree Search (ULTS) which places a prior belief over any pre-trained LLM’s softmax outputs (which are discrete distributions) and computes the implied posterior beliefs over the optimal likelihood values. We proceed by introducing and discussing the modeling assumptions in Section 3.1 and show how to derive approximate beliefs over the optimal values in the search tree in Section 3.2. The samples of the posterior beliefs are then used to make decisions about which subtree to expand next and when to stop the search in Section 3.4. Figure 2 gives an overview of our probabilistic model.
@@ 3.1 Prior beliefs over LLMs softmax outputs
A straightforward belief one can consider is the Dirichlet distribution. For tractability, we assume that the LLM’s softmax probabilities are iid. draws from a symmetric Dirichlet distribution with concentration parameter $\alpha>0$, i.e., $p(\boldsymbol{c}_{i})=\mathrm{Dir}(\alpha)$. Thus, $\alpha$ controls how concentrated the sampled probability vectors are. In the context of LLMs, for small $\alpha$, the LLM would typically strongly favor a few tokens, whereas for large $\alpha$ the discrete distribution would closely resemble a uniform distribution over the tokens. The symmetry of the prior implies in our context that we do not have a preference for particular tokens a priori.
Another belief that we can consider is an empirical prior over $\boldsymbol{c}_{i}$. Let $\smash{\{\boldsymbol{x}_{n}\}_{n=1}^{N}}$ be samples of contexts, e.g. a subset of the training/validation data. Given an LLM $f$, we can then obtain the set $\{(\boldsymbol{x}_{i},x_{i1},\dots,x_{id})\}$ of $d$-step completions of $\boldsymbol{x}_{i}$’s through the LLM (e.g. through a greedy decoding). We can then collect samples of the Categorical distributions from this generation process: $\mathcal{C}=\smash{\coprod_{n=1}^{N}}\{\boldsymbol{c}_{n1},\dots,\boldsymbol{c}_{nd}\}$. Then, instead of sampling from $\mathrm{Dir}(\alpha)$, we can sample from $p(\boldsymbol{c}_{i})=\mathrm{Unif}(\mathcal{C})$. This prior is more flexible than the Dirichlet prior since no symmetric nor unimodality assumptions are made, at the cost of more compute and memory overhead. Note however that all these priors are precomputed and can be reused across subsequent decoding runs so they incur a fixed $\mathcal{O}(1)$ cost.
3.2 Prior beliefs over the optimal values
@@
Having picked a prior over the LLM’s softmax outputs, we can derive the implied priors over the optimal values $v_{x_{i}}$ for each node $x_{i}$ in the tree [11]. From the definition of an optimal value (Section 2), it factorizes as the product $c_{\boldsymbol{x}_{0}\rightarrow x_{i}}$ of the transition probabilities from the root node $x_{0}$ to $x_{i}$ and a remaining term which we refer to as $\Delta_{i}$, i.e., we have $v_{x_{i}}=c_{\boldsymbol{x}_{0}\rightarrow x_{i}}\cdot\Delta_{i}$. Intuitively, the term $\Delta_{i}$, quantifies the likelihood that we get in the remaining steps from $x_{i}$ to a leaf node when we take all remaining decisions optimally. It can be defined by the following recurrence relation:
@@
Due to the iid. assumption above (Section 3.1), we have the joint distribution $p(c_{\boldsymbol{x}_{0}\rightarrow x_{i}},\Delta_{i})=p(c_{\boldsymbol{x}_{0}\rightarrow x_{i}})\,p(\Delta_{i})$. Note that, all these quantities are not analytically available. Thankfully, sampling from the posterior belief $p(v_{x_{i}}\mid c_{\boldsymbol{x}_{0}\to x_{i}})$ is easy if we are able to sample from $p(\Delta_{i})$.
We now derive an approximate sampling scheme for $p(\Delta_{i})$ (Algorithm 1). It follows the one used in [11, 8] with Gaussian priors. We recursively approximate the prior distribution of the $\Delta_{i}$’s at level $l$ with Beta distributions $\mathcal{B}_{l}(\Delta_{i})$. In a bottom-up approach, we generate samples
$\left\{\max_{j}c_{nj}\mid c_{n}\sim p(\boldsymbol{c})\right\}_{n=1}^{N}$
@@
for a $\Delta_{i}$ at level $l=d-1$. Using these samples, we empirically fit the parameters of the Beta distribution $\mathcal{B}_{d-1}(\Delta_{i})$ via maximum likelihood [1]. The distributions of the $\Delta_{i}$ are the same for all nodes on the same level due to the iid. assumption, so this has to be done only once. Note that we need this approximation since the distribution of the maximum $\max_{j}c_{ij}$ has no known analytic solution. We then continue, by recursively sampling sets of the form (one per level)
$\left\{\max_{j}c_{nj}\cdot\Delta_{j}\mid c_{n}\sim p(\boldsymbol{c}),\Delta_{j}\sim\mathcal{B}_{l+1}(\Delta)\right\}_{n=1}^{N}$
for a $\Delta_{i}$ of the level $l$ and using it to fit the parameters of $\mathcal{B}_{l}(\Delta_{i})$. The time complexity is $\mathcal{O}(d\cdot b\cdot N)$ for computing the approximations, i.e., it is linear in the depth and width of the tree. We emphasize that they can be computed before the search and reused across sentences, i.e., these costs are irrelevant to the decoding process itself.
3.3 Posterior beliefs over the optimal values

### v2

@@ Abstract
Tree search algorithms such as greedy and beam search are the standard when it comes to finding sequences of maximum likelihood in the decoding processes of large language models (LLMs). However, they are myopic since they do not take the complete root-to-leaf path into account. Moreover, they are agnostic to prior knowledge available about the process: For example, it does not consider that the objective being maximized is a probability and thereby has specific properties like being bound in the unit interval. Taking a probabilistic approach, we define prior beliefs over LLMs’ transition probabilities and obtain posterior beliefs over the most promising paths in each iteration. These beliefs are useful for defining a sample-based, non-myopic acquisition function that allows for a more data-efficient exploration scheme than standard search algorithms on LLMs. Crucially, unlike expensive simulation-based non-myopic methods like the Monte Carlo tree search, our method only requires samples from the beliefs. Our formulation thus views LLM decoding as Bayesian optimization on trees. We discuss how to select the prior and the acquisition function, and demonstrate in experiments with various LLMs that our method achieves higher efficiency than recent baselines: Our method achieves the same or a higher likelihood while expanding fewer nodes.
@@ 1 Introduction
The decoding process of an autoregressive large language model (LLM) can be seen as finding an optimal path in a search tree. The number of such paths is exponential, often exceeding the computational budget required to examine them all. This inevitably leads to computational uncertainty (Hennig et al., 2022): an uncertainty that could be fully resolved if enough compute was available to examine all paths, but in practice is present due to the limited resources. Standard algorithms for LLM decoding, such as beam search (Koehn et al., 2003), completely ignore this uncertainty.
@@
We posit that quantifying this uncertainty can be beneficial to ensure better explorations on LLMs’ search trees. To this end, we incorporate computational uncertainty into the search process to guide it in a non-myopic fashion (i.e., accounting for our belief about the values of future nodes) and importantly, in a more data-efficient manner, akin to Bayesian optimization methods (BO, Kushner, 1964; Močkus, 1975; Garnett, 2023). BO methods are recognized for their data efficiency, not merely because they quantify uncertainty, but because they exploit the structural characteristics within that uncertainty. E.g., in continuous optimization problems, prior knowledge, such as the smoothness of a function, is often available through Gaussian processes (Rasmussen & Williams, 2005) or (Bayesian) neural networks (Hernández-Lobato et al., 2017; Kristiadi et al., 2023; 2024).
@@
Figure 1: The intuition of our method—faded color represents unexplored subtrees. We sequentially expand a node of the LLM search tree based on the posterior belief. The latter is obtained by conditioning the unknown (the optimal total likelihood $\max c_{x_{0}\to x_{i}\to x_{n}}$ from the root $\boldsymbol{x}_{0}$ to a leaf $x_{n}$ when continuing down a particular node $x_{i}$) on the current observations (the total likelihood $c_{\boldsymbol{x}_{0}\to x_{i}}$ up until that node). This belief is induced by a simple prior belief on the LLM softmax probabilities $\boldsymbol{c}$. The samples of the posterior belief are used to decide which child node should be expanded next.
@@
Unlike standard BO, however, LLM decoding is a structured, discrete optimization problem—its search space is the set of all root-leaf paths in a tree. Moreover, commonly, the values or rewards associated with each node are probabilities, bounded between 00 and $1$. Hence, uncertainty quantification here is not as well-studied as in standard BO. We will, therefore, assume that rewards at a node of the search tree are components of a discrete distribution. The characteristic property we aim to exploit is the concentration strength: whether the Categorical distributions are all highly concentrated at a few realistic options, or if some of them are nearly uniform, making any individual option drastically less likely to be the optimal one. Intuitively, one would expect this to have a strong influence on the number of paths that need to be considered. For instance, when the distribution is concentrated, it is less likely under our belief that other paths will overtake later on and one can be more greedy. Meanwhile, when the distribution has high entropy, the uncertainty of our belief about “which next token is best” is higher and thereby requires more exploration and computational budget.
In this work, we propose a probabilistic framework that captures this aspect of the search space, which can help to decide which paths should be pursued and which can be ignored based on the posterior belief (Fig. 1). We do so by using a non-myopic acquisition function based on the samples of such a belief, which can be seen as a generalization of Thompson sampling (Thompson, 1933). Crucially, these samples are cheap, unlike other sample-based tree-search methods like Monte Carlo tree search. Moreover, we show that our acquisition function can easily be extended, e.g., to prefer sequences with higher diversity and less text degeneration (Holtzman et al., 2019).
@@
Experiments on real-world text-generation benchmarks with various LLMs suggest that our method, called Uncertainty-guided Likelihood-Tree Search (ULTS), finds sentences with higher rewards than recent baselines with significantly fewer expensive node expansions. Moreover, ULTS only adds a small runtime overhead relative to the forward passes of the LLM. In summary:
@@
    We propose ULTS: a probabilistic decision-making framework on LLM search trees, where a prior belief is put on LLM softmax outputs. We show how to easily sample from the implied posterior over optimal values, and how to use these samples to make decisions when searching the tree. Our work thus bridges natural language processing with Bayesian optimization.
@@
    We demonstrate the efficiency and extensibility of ULTS in decoding recent LLMs.
@@
    We open-source an implementation compatible with transformers (Wolf et al., 2020) that is available at https://github.com/JuliaGrosse/ults.
@@ 2.1 Decoding large language models
Let ${\mathcal{A}}=\{a_{1},\dots,a_{b}\}$ be a vocabulary—a set of natural-language tokens. A large language model (LLM) is a neural network that utilizes the attention mechanism (Vaswani et al., 2017), taking a context, an ordered sequence of tokens $\boldsymbol{x}_{0}=(x_{01},\dots,x_{0k})$—each component takes values in the set ${\mathcal{A}}$—and outputs a distribution over ${\mathcal{A}}$, i.e. $\boldsymbol{c}=(c_{1},\dots,c_{b})$ where each $c_{i}\geq 0$ and $\smash{\sum_{i=1}^{b}}c_{i}=1$.
@@
LLM decoding can be seen as a tree-search problem (Fig. 1). The context $\boldsymbol{x}_{0}$ is the root of the tree and at each step $i=1,\dots,d$, we are given a choice of $b$ many tokens to pick. This process is done recursively until termination; either when a specified depth $d$ has been reached or when a special token like “<EOS>” is selected. This means, the LLM search tree is an exponentially large tree (w.r.t. the number of tokens $b$ in the vocabulary) with depth $d$—the number of paths $\boldsymbol{x}_{0}\to x_{d}$ from the root to leaves is $b^{d}$. Considering the fact that $b$ ranges from around $32$k to $256$k (Radford et al., 2019; Chowdhery et al., 2023), generating a sequence of tokens requires search over an intractably large space.
@@
To address this problem, one can use a cheaper but heuristic way to explore this tree. The simplest way to do so is by sampling a token from the distribution $\boldsymbol{c}$ at each level of the tree (Holtzman et al., 2019). In some domains like machine translation and summarization, however, one often wants to find the sequence that maximizes the total/joint likelihood  [^1] Implementation wise, this is usually done in the logarithmic space. $c_{\boldsymbol{x}_{0}\to x_{d}}:=\smash{\prod_{i=1}^{d}}c_{(i-1)i}$ associated with each possible sequence $(\boldsymbol{x}_{0},x_{1},\dots,x_{d})$ (Wiher et al., 2022). This is an optimization-on-tree problem—heuristic optimization algorithms like greedy and beam search along with their variants (Vijayakumar et al., 2018; Meister et al., 2021; Freitag & Al-Onaizan, 2017, etc.) are the standard.
@@
One can also view the decoding process as follows. On each of the tree’s node $x_{i}$, we can define an optimal value $v_{x_{i}}$ associated with it: If $x_{d}$ is a leaf node, then $v_{x_{d}}$ corresponds to the total likelihood $c_{\boldsymbol{x}_{0}\rightarrow x_{d}}$ from the root until that leaf node. Meanwhile, for inner nodes $x_{i}$ it is defined recursively in a bottom-up fashion as the maximum of the children’s optimal values $\max_{x_{c}\in\mathrm{children}{(}x_{i})}{v_{x_{c}}}$. Intuitiveley, $v_{x_{i}}$ tells us “what is the total likelihood we would get if we continue descending the tree through the node $x_{i}$ in an optimal manner”. The optimization-on-tree problem can then be recast as finding a path $\boldsymbol{x}_{0}\to x_{d}$ that corresponds to the optimal value $v^{*}$ of the root. While for all $x_{i}$, the quantity $v_{x_{i}}$ cannot be feasibly computed, they are useful for our probabilistic treatment of LLM decoding in Section 3.
2.2 Probabilistic decision-making on trees
Bayesian optimization (Kushner, 1964; Močkus, 1975) approaches black-box optimization as a probabilistic sequential decision-making problem. This is done by first modeling a belief about the unknown given the current observation. This belief is then used to make decision, by computing an acquisition function which scores possible candidates. The candidate that has the highest acquisition value is then selected and the process is repeated until termination.
@@ 2.2 Probabilistic decision-making on trees
In the context of tree-structured problem, the approach to perform probabilistic decision-making is to perform probabilistic modeling over the unknown values $v_{x_{i}}$’s and use the posterior beliefs to make decision about which path should be followed next. This kind of optimization has been done for game (e.g., Go) trees and directed-acyclic graphs (Hennig et al., 2010; Grosse et al., 2021). They assume that the optimal value $v_{x_{i}}$ is decomposed into a latent score $g_{i}$ for the utility of the node if all remaining steps are taken randomly and additional increment $\Delta_{i}$ to quantify how much better this score can get when all remaining steps are taken optimally instead of randomly, i.e. $v_{x_{i}}=g_{i}+\Delta_{i}$. They assume a Gaussian process with the Brownian motion kernel as their prior belief.
@@ 3 ULTS: Uncertainty-guided Likelihood-Tree Search
Here, we discuss Uncertainty-guided Likelihood-Tree Search (ULTS) which places a prior belief over any pre-trained LLM’s softmax outputs and computes the implied posterior beliefs over the optimal likelihood values. We introduce and discuss the modeling assumptions in Section 3.1 and show how to derive approximate beliefs over the optimal values in the search tree in Section 3.2. The samples of the posterior beliefs are then used to make decisions about which subtree to expand next and when to stop the search (Section 3.4). Figure 2 gives an overview of our probabilistic model.
@@ 3.1 Prior beliefs over LLMs softmax outputs
A straightforward belief one can consider is the Dirichlet distribution. For tractability, we assume that the LLM’s softmax probabilities are iid. draws from a symmetric Dirichlet distribution with parameter $\alpha>0$, i.e., $p(\boldsymbol{c}_{i})=\mathrm{Dir}(\alpha)$. Thus, $\alpha$ controls how concentrated the sampled probability vectors are. In the context of LLMs, for small $\alpha$, the LLM would typically strongly favor a few tokens, whereas for large $\alpha$ the discrete distribution would closely resemble a uniform distribution over the tokens. The symmetry of the prior implies in our context that we do not have a preference for particular tokens a priori.
Another belief that we can consider is an empirical prior over $\boldsymbol{c}_{i}$. Let $\smash{\{\boldsymbol{x}_{n}\}_{n=1}^{N}}$ be samples of contexts, e.g. a subset of the training/validation data. Given an LLM $f$, we can then obtain the set $\{(\boldsymbol{x}_{i},x_{i1},\dots,x_{id})\}$ of $d$-step completions of $\boldsymbol{x}_{i}$’s through the LLM (e.g. through a greedy decoding). We can then collect samples of the Categorical distributions from this generation process: $\mathcal{C}=\smash{\coprod_{n=1}^{N}}\{\boldsymbol{c}_{n1},\dots,\boldsymbol{c}_{nd}\}$. Then, instead of sampling from $\mathrm{Dir}(\alpha)$, we can sample from $p(\boldsymbol{c}_{i})=\mathrm{Unif}(\mathcal{C})$. This prior is more flexible than the Dirichlet prior since no symmetric nor unimodality assumptions are made at the cost of more compute and memory overhead. Note, however, that all these priors are precomputed and can be reused across subsequent decoding runs, so they incur a fixed $\mathcal{O}(1)$ cost.
3.2 Prior beliefs over optimal values
@@
Having picked a prior over the LLM’s softmax outputs, we can derive the implied priors over the optimal values $v_{x_{i}}$ for each node $x_{i}$ in the tree (Hennig et al., 2010). From the definition of an optimal value (Section 2), it factorizes as the product $c_{\boldsymbol{x}_{0}\rightarrow x_{i}}$ of the transition probabilities from the root node $x_{0}$ to $x_{i}$ and a remaining term which we refer to as $\Delta_{i}$, i.e., we have $v_{x_{i}}=c_{\boldsymbol{x}_{0}\rightarrow x_{i}}\cdot\Delta_{i}$. Intuitively, the term $\Delta_{i}$, quantifies the likelihood that we get in the remaining steps from $x_{i}$ to a leaf node when we take all remaining decisions optimally. It can be defined by the following recurrence relation:
@@
Due to the iid. assumption above (Section 3.1), we have the joint distribution $p(c_{\boldsymbol{x}_{0}\rightarrow x_{i}},\Delta_{i})=p(c_{\boldsymbol{x}_{0}\rightarrow x_{i}})\,p(\Delta_{i})$. Note that all these quantities are not analytically available. Thankfully, sampling from the posterior belief $p(v_{x_{i}}\mid c_{\boldsymbol{x}_{0}\to x_{i}})$ is easy if we are able to sample from $p(\Delta_{i})$ (Section 3.3).
@@
Let us, therefore, derive an approximate sampling scheme for $p(\Delta_{i})$ (Alg. 1). It follows Hennig et al. (2010); Grosse et al. (2021) who used Gaussian priors. We recursively approximate the prior distribution of the $\Delta_{i}$’s at level $l$ with Beta distributions $\mathcal{B}_{l}(\Delta_{i})$. In a bottom-up approach, we generate samples $\{\max_{j}c_{nj}\mid c_{n}\sim p(\boldsymbol{c})\}_{n=1}^{N}$ for a $\Delta_{i}$ at level $l=d-1$. Using these samples, we empirically fit the parameters of the Beta distribution $\mathcal{B}_{d-1}(\Delta_{i})$ via maximum likelihood (AbouRizk et al., 1994). The distributions of the $\Delta_{i}$ are the same for all nodes on the same level due to the iid. assumption, so this has to be done only once. Note that we need this approximation since the distribution of the maximum $\max_{j}c_{ij}$ has no known analytic solution. We then continue by recursively sampling sets of the form (one per level) $\{\max_{j}c_{nj}\cdot\Delta_{j}\mid c_{n}\sim p(\boldsymbol{c}),\Delta_{j}\sim\mathcal{B}_{l+1}(\Delta)\}_{n=1}^{N}$ for a $\Delta_{i}$ of the level $l$ and using it to fit the parameters of $\mathcal{B}_{l}(\Delta_{i})$. The time complexity is $\mathcal{O}(d\cdot b\cdot N)$ for computing the approximations, i.e., it is linear in the depth and width of the tree. We emphasize that they can be pre-computed before the search and reused across different decoding runs.
3.3 Posterior beliefs over optimal values